# Can the 3 neutrino masses really be found using SN 1987A data?

### v4

@@ Abstract
Neutrino masses remain a significant unsolved problem in physics and their nonzero value proves the Standard Model is incomplete. Currently, the values of the three masses only have upper limits from cosmology and experiments like KATRIN. This paper shows that the SN 1987A neutrino data can remarkably yield values for the three neutrino masses, and not merely upper limits. Although this seemingly preposterous idea was suggested a dozen years ago by the author, here it is demonstrated in a much more convincing manner with many new elements, including a stronger statistical treatment, and most importantly, a thorough explanation of why the method used to find the three masses from supernova SN 1987A neutrino data really works. The key to finding the three neutrino masses is realizing why three normally accepted assumptions are unjustified, The three rejected assumptions are:(a) the 5-hr early LSD (Mont Blanc) neutrinos are unrelated to SN 1987A, (b) any neutrino masses $m_{k}>1eV/c^{2}$ cannot be reconciled with upper limits on the “effective mass” from KATRIN and other data, and (c) the spread in neutrino emission times from SN 1987A data is too great for the method to work. A particularly crucial piece of evidence supporting the claim made in the paper’s title involves a recent negative KATRIN result finding an absence of sterile neutrinos.
Keywords: SN 1987A, Neutrino, supernova, violation, tachyon, KATRIN, sterile neutrino
@@ I Introduction
Thus, it is usually assumed that the spread in neutrino emission times for SN 1987A is large enough to disguise any tiny differences in neutrino travel times from which the neutrino mass could be found, Here (and elsewhere) we make no such assumption, and instead use the SN 1987A data themselves to see what they reveal about whether the masses can be found. As a result, the author rediscovered what Ramanath Cowsik had found in 1988 about the neutrino masses being possibly revealed in the SN 1987A data. (Ehrlich2011, ) As we shall see, the data suggests $m^{2}$ values for the three neutrino masses, all of which are much larger than thought possible, and one of which is a tachyon, i.e., $m^{2}<0.$ (Bilaniuk1962, ), While this paper is not the first time the author has made this claim, there is much new here, including (1) a much improved statistical analysis, (2) a demonstration that finding individual neutrino masses is possible, given the best current neutrino emission model of SN 1987A and (3) an invaluable clue from the KATRIN experiment’s reported failure to observe sterile neutrinos. This absence dramatically strengthens the case that the three neutrino masses really can be found from the SN 1987A data, and it suggests a way KATRIN can redo its analysis that may yield a confirmation of the preposterous-sounding claim. (KATRIN2025, ). The paper also suggests a way that short baseline $\nu_{e}$ disappearance experiments can supply such a confirmation.
@@ IV.8 Consistency with cosmological data?
so, the appropriate constraint on R from fitting cosmological data is from Eq. 6 a constraint on the flavor state sum, and not the mass state sum, as sometimes assumed. Thus, even though the three mass state $m^{2}$ values from the SN 1987A fit are all very large, as long as one of them has $m^{2}<0,$ the three flavor state effective masses can all be very tiny, which appears to be the case.
@@ IV.9 Consistency with direct mass experiments?
To test the validity of the main claim of this paper (that the three mass fit of SN 1987A data yields the three true masses), KATRIN could see where on the $\alpha_{1}$ and $\alpha_{2}$ plane good fits to the observed spectrum are found. With the three masses set to their specified SN 1987A values, KATRIN should observe a small concentrated area where good fits are found surrounded by a large area of poor fits. They could also do the same using other masses, where no good fits should be found. The best fit on the $\alpha_{1}$ and $\alpha_{2}$ plane should yield the effective mass from $m^{2}=\sum\alpha_{j}m^{2}_{j},$ and it should be a very small negative value.
Finally, for the functional form of the spectrum for the tachyonic third mass, $m_{3}^{2},$ they could just use the usual beta decay formula, and simply substitute a negative $m^{2}$ in it. While Rembielinski, Caban and Ciborowski have shown that their causal tachyonic neutrino model results in a more complicated formula – see Eqs. 93-94 in ref. (Rembielinski2021, ), however, as shown in their Fig. 1 the simple approach gives a result that is virtually indistinguishable from their full description.
@@ IV.12 Multiple sterile neutrinos in oscillation data?
Anatolii Serebrov and his Neutrino 4 team, who first detected reactor antineutrino oscillations in the 6-12 meter range, has advocated the need for two more sterile neutrinos in addition to $m_{1}^{2}=2.70^{2}eV^{2}/c^{4}$ based on theoretical grounds (Serebrov2024, ) Assuming a second sterile neutrino of mass $m_{2}^{2}=22.4eV^{2}/c^{2},$ the oscillations would presumably show up in a range roughly 80 times shorter. If such oscillations are detectable one approach to find them would involve setting up a grid of values for the two masses $m_{1},m_{2}$ and in each case use the data to find thelikelihood probability of the fit to the data for that pair of mass values. One could then see where on the plane that likelihood maximizes. If there is a peak right at a small region of the grid near $(m^{2}_{1},m_{2}^{2})=(2.70^{2},22.4^{2})eV^{2}/c^{2},$ that would constitute strong support for the idea that the $m^{2}$ masses of electron neutrinos consist of the values deduced from the SN 1987A data. The most promising experiments to look at might be the four identified by Serebrov in ref. (Serebrov2024, ). These already have shown evidence of the $2.70eV/c^{2}$ mass, and the question is do they now also show evidence of a $22.4eV/c^{2}$ neutrino. It would seem highly unlikely that a signature of any oscillation due to the $m_{3}^{2}=-200,000eV^{2}/c^{4}$ neutrino could be found in a similar manner, but it should be detectable in KATRIN as already discussed.
@@ V Summary and the ultimate test
This paper has made six claims that many. readers may understandably consider very difficult to believe – see Table 1. Some of the claims will likely not be accepted as true without the strongest corroborating evidence. After KATRIN does the analysis suggested in section IV.10, and it were to find evidence for the three active neutrino masses deduced from the fit to SN 1987 data with little uncertainty in their values, that might be enough. This might also be the case if oscillation experiments were to find such evidence doing the analysis mentioned in the previous section.
1. The LSD 5-hr early burst was connected to SN 1987A
@@ 4. One of the 3 $m^{2}_{j}$ from SN 1987A data is a $m^{2}<0$ tachyon
6. Existing data likely can confirm or reject all the above
@@
However, the most convincing test would be a new galactic supernova. Given the much larger neutrino detectors today compared to those in 1987, a new galactic supernova might yield thousands of neutrinos reaching the detectors. This 88 year-old does not have a great chance of seeing it, but let’s be an optimist. When supernova SN 2026(!) occurs at a distance that is a percentage p of the distance of SN 1987A, and we were to find an LSD-like near-monochromatic burst of a thousand neutrinos occurring earlier than the main neutrino burst by a time interval of $484\times p$ minutes, there would be little doubt that tachyons exist, as well as the other five claims. Even if the next galactic supernova does not occur in 2026, and I don’t get to see it, it is comforting to realize the evidence, in the form of thousands of neutrinos are now on their way to Earthly detectors. These neutrinos will definitely either confirm or reject the six claims made. Moreover, given the known frequency of supernovae, we know that those neutrinos are now more than $99.99\%$ of their way here from some star in our galaxy. No one knows which star it will be, but Betelgeuse, a massive red supergiant in Orion, is expected to end its life in a spectacular supernova explosion soon. Soon in astronomical terms means potentially within the next 100,000 years (some say under 300 years (Saio2023, )). Were Betelgeuse to become a supernova, it would become incredibly bright, potentially as bright as the full moon. Go Betelgeuse!

### v5

@@ Abstract
Neutrino masses remain a significant unsolved problem in physics and their nonzero value proves the Standard Model is incomplete. Currently, the values of the three masses only have upper limits from cosmology and experiments like KATRIN. This paper shows that the SN 1987A neutrino data can remarkably yield values for the three neutrino masses, and not merely upper limits. Although this seemingly preposterous idea was suggested a dozen years ago by the author, here it is demonstrated in a much more convincing manner with many new elements, including a stronger statistical treatment, and most importantly, a thorough explanation of why the method used to find the three masses from supernova SN 1987A neutrino data really works. The key to finding the three neutrino masses is realizing why three normally accepted assumptions are unjustified, The three rejected assumptions are:(a) the 5-hr early LSD (Mont Blanc) neutrinos are unrelated to SN 1987A, (b) any neutrino masses $m_{k}>1eV/c^{2}$ cannot be reconciled with upper limits on the “effective mass” from KATRIN and other data, and (c) the spread in neutrino emission times from SN 1987A data is too great for the method to work. A particularly crucial piece of evidence supporting the claim made in the paper’s title involves a recent negative KATRIN result finding an absence of sterile neutrinos. This absence of a sterile neutrino signal leads to two tests of the claim based on existing data: one for KATRIN and one for IceCube.
Keywords: SN 1987A, Neutrino, supernova, tachyon, KATRIN, IceCube, sterile neutrino
@@ I Introduction
Thus, it is usually assumed that the spread in neutrino emission times for SN 1987A is large enough to disguise any tiny differences in neutrino travel times from which the neutrino mass could be found, Here (and elsewhere) we make no such assumption, and instead use the SN 1987A data themselves to see what they reveal about whether the masses can be found. As a result, the author rediscovered what Ramanath Cowsik had found in 1988 about the neutrino masses being possibly revealed in the SN 1987A data. (Ehrlich2011, ) As we shall see, the data suggests $m^{2}$ values for the three neutrino masses, all of which are much larger than thought possible, and one of which is a tachyon, i.e., $m^{2}<0.$ (Bilaniuk1962, ), While this paper is not the first time the author has made this claim, there is much new here, including (1) a much improved statistical analysis, (2) a demonstration that finding individual neutrino masses is possible, given the best current neutrino emission model of SN 1987A and (3) an invaluable clue from the KATRIN experiment’s reported failure to observe sterile neutrinos. This absence dramatically strengthens the case that the three neutrino masses really can be found from the SN 1987A data, and it suggests a way KATRIN can redo its analysis that may yield a confirmation of the preposterous-sounding claim. (KATRIN2025, ). The paper also suggests a way that oscillation experiments and especially the IceCube experiment might supply such a confirmation.
@@ IV.8 Consistency with cosmological data?
so, the appropriate constraint on R from fitting cosmological data is from Eq. 6 a constraint on the flavor state sum, and not the mass state sum, as sometimes assumed. Thus, even though the three mass state $m^{2}$ values from the SN 1987A fit are all very large, as long as one of them has $m^{2}<0,$ the three flavor state effective masses can all be very tiny, which appears to be the case. Conversely, for the neutrino mass sum to be a sum over mass states, there would need to be a process that created neutrinos in equal numbers according to their mass state, and there is no such process.
@@ IV.9 Consistency with direct mass experiments?
To test the validity of the main claim of this paper (that the three mass fit of SN 1987A data yields the three true masses), KATRIN could see where on the $\alpha_{1}$ and $\alpha_{2}$ plane good fits to the observed spectrum are found. With the three masses set to their specified SN 1987A values, KATRIN should ideally observe a small concentrated area where good fits are found surrounded by a large area of poor fits. They could also do the same using other masses, where no good fits should be found. The best fit on the $\alpha_{1}$ and $\alpha_{2}$ plane should yield the effective mass from $m^{2}=\sum\alpha_{j}m^{2}_{j},$ and it should be a very small negative value.
Finally, for the functional form of the spectrum for the tachyonic third mass, $m_{3}^{2},$ KATRIN could just use the usual beta decay formula, and simply substitute a negative $m^{2}$ in it. While Rembielinski, Caban and Ciborowski have shown that their causal tachyonic neutrino model results in a more complicated formula – see Eqs. 93-94 in ref. (Rembielinski2021, ), as shown in their Fig. 1 the simple approach gives a result that is virtually indistinguishable from their full description.
@@ IV.12 Multiple sterile neutrinos in oscillation data?
Anatolii Serebrov and his Neutrino 4 team, have argued on cosmological grounds for there being two heavier sterile neutrinos in addition to the one found at $m_{1}^{2}=2.70^{2}eV^{2}/c^{4},$ which of course was also earlier inferred from the SN 1987A data analysis. Assuming a second sterile neutrino of mass $m_{2}^{2}=22.4eV^{2}/c^{2},$ as mandated by that data, the oscillation wavelength would be 69 times shorter than the $m_{1}$ sterile neutrinos. It is highly unlikely that such short wavelength oscillations are detectable in the data of Neutrino 4, but if they were one approach to find them would involve setting up a grid of values for the two masses $m_{1},m_{2}$ and in each case use the data to find the likelihood probability of the fit to the data for that pair of mass values. One could then see where on the plane that likelihood maximizes. If there is a peak right at a small region of the grid near $(m^{2}_{1},m_{2}^{2})=(2.70^{2},22.4^{2})eV^{2}/c^{2},$ that would constitute strong support for the idea that the $m^{2}$ masses of electron neutrinos consist of the values deduced from the SN 1987A data. You might think that the most promising experiments to look at would be the four data sets identified by Serebrov in ref. (Serebrov2024, ), which already have shown evidence of the $2.70eV/c^{2}$ mass at a cumulative $5.8\sigma$. Unfortunately, however, the very short wavelength of $22.4^{2}eV^{2}/c^{2}$ neutrinos almost certainly make these oscillations undetectable in existing short baseline reactor data. There may be over 20 projects in the works exploring the ultra-short baseline region ($L<10$ meters) in reactor experiments, (Serebrov2021, ) but no experiments known to the author could detect stable sterile neutrinos of such a high mass, with the exception of IceCube.
In fact, IceCube has sought evidence for $\nu_{\mu}$ and $\bar{\nu_{\mu}}$} disappearance oscillations in seven years of data. (Abbasi2024, ) Even though they reported a negative result their best fit did show a hint of a value consistent with the $m_{1}=2.70eV/c^{2}$ mass whuch was strongly preferred over the no oscillation solution by 10.7 to one. Furthermore, the IceCube energy range (up to $E=100TeV$) would make them sensitive to sterile mass values as high as $m_{2}^{2}=22.4^{2}eV^{2}/c^{4}.$ Thus, it might make sense for them to rexamine their data for evidence of a double oscillation for those two specific masses. Making such a search even more promissing, IceCube has already begun to explore the region of $\Delta m^{2}>10eV^{2}/c^{4},$ and found a preference for $3+1$ oscillations. (Biennow2018, ) A search using the method outlined for the Neutrino-4 data might prove particularly promising. Although the SN 1987A data analysis specified three neutrino masses it seems highly unlikely that the $m_{3}^{2}=-200,000eV^{2}/c^{4}$ neutrino would leave any detectable signal in IceCube or any other oscillation data (unlike KATRIN where it should be detectable).
@@ V Summary and the ultimate test
This paper has made six claims that many readers may understandably consider very difficult to believe – see Table 1. Some of the claims will likely not be accepted as true without the strongest corroborating evidence. After KATRIN does the analysis suggested in section IV.10, and it were to find evidence for the three active neutrino masses deduced from the fit to SN 1987 data, that might be enough. This might also be the case if oscillation experiments were to find such evidence doing the analysis mentioned in the previous section.
1. The LSD burst was due to SN 1987A but no 2nd bang
@@ 4. One of the 3 $m^{2}_{j}$ from SN 1987A data is a $m^{2}<0$ tachyon
6. KATRIN, IceCube or other data likely can confirm #5
@@
However, the most convincing test would be a new galactic supernova. Given the much larger neutrino detectors today compared to those in 1987, a new galactic supernova might yield thousands of neutrinos reaching the detectors. This 88 year-old does not have a great chance of seeing it, but let’s be an optimist. When supernova SN 2026(!) occurs at a distance that is a percentage p of the distance of SN 1987A, and we were to find an LSD-like near-monochromatic burst of a thousand neutrinos occurring earlier than the main neutrino burst by a time interval of $484\times p$ minutes, there would be little doubt that tachyons exist, as well as the other five claims. Even if the next galactic supernova does not occur in 2026, and I don’t get to see it, it is comforting to realize the evidence, in the form of thousands of neutrinos are now on their way to Earthly detectors.
These neutrinos will definitely either confirm or reject the six claims made. Moreover, given the known frequency of supernovae, we know that those neutrinos are now more than $99.99\%$ of their way here from some star in our galaxy. No one knows which star it will be, but Betelgeuse, a massive red supergiant in Orion, is expected to end its life in a spectacular supernova explosion soon. Soon in astronomical terms means potentially within the next 100,000 years (some say under 300 years (Saio2023, )). Were Betelgeuse to become a supernova, it would become incredibly bright, potentially as bright as the full moon. Go Betelgeuse!